Find standard deviation for the following returns:

Stock XYZ

2011 -18.00%

2012 30.00

2013 15.00

2014 -2%

2015 27.00 %
Answer #1

Given Returns: 18 %, 30%, 15 %, - 2 % and 27 %, Number of Years = 5

Mean Return = (18 + 30 + 15 - 2 + 27) / 5 = 17.6 %

Mean of Squared Deviations = [(18-17.6)^(2) + (30 - 17.6)^(2) + (15 - 17.6)^(2) + (-2 - 17.6)^(2) + (27 - 17.6)^(2)] / 5 = 633.2 / 5 = 126.64

Standard Deviation = Square Root of Mean of Squared Deviations = (126.64)^(1/2) = 11.2534 % ~ 11.25 %